A convex spherical mirror has a radius of curvatur
Answer #1

a. focal length = f= R/2 = 46/2= 23 cm

1/f= 1/do+1/di

so, 1/23=-1/41+ 1/di

so, di= 14.73 cm

magnification =14.73/41=0.359

b. focal length = f= R/2 = 46/2= 23 cm

1/f= 1/do+1/di

so, 1/23=-1/55+ 1/di

so, di= 16.21 cm

magnification =16.21/55=0.294

c.image is upright

image is upright